# Notes on Correlation Stress Tests

### v1

@@ 1. Introduction
This note is organised as follows: the next section introduces certain well-known features of the Fisher–Rao manifold. Section 3 discusses the meaning of covariance and correlations stress tests and Section 4 discusses the interpretation of correlation stress tests as geodesics in the Fisher–Rao manifold. Section 5 proposes an explicit parametrisation of the correlation stress tests which is the exponential map of the tangent space at the un-stressed covariance. Section 6 discusses the plausibility of the stress tests, while sections 7 through 9 discuss the stress tests that result from specific choices of “directions” in which the correlation matrix is stressed. In Section 10 we show that a correlation stress test which affects all correlations by the same amount results in the same lowest-order functional form as the estimation bias of eigenvalues of the covariance matrix due to finite amount of data. Section 11 reflects on certain features of the stress tests defined by a Lax-pair evolution that would make them less suitable as a model for correlation stress test. Section 12 illustrates the main concepts using a stylised example of bonds, equities and commodities. The results are summarised in Section 13.
@@ 2. The Fisher–Rao manifold
[…]
@@
[…]
More specifically, consider the family of multivariate Gaussian distributions with vector of means $\mu$ and covariance $\Sigma$, and assume that both $\mu$ and $\Sigma$ are differentiable functions of some arbitrary parameters $\theta_{i}$, $i=1,\dots,k$ The Fisher–Rao metric reduces to [12, 13]:
@@ 3. Covariance and correlation stress tests
A geometric interpretation provides some further guidance: the determinant of the covariance matrix is, for a multivariate normal distribution, the volume of an ellipsoid of constant probability density enveloping a given probability mass. The directions of the semi-axes of this ellipsoid represent the uncorrelated directions (principal components), while the length of a semi-axis represents the standard deviation of the principal component. A change in the covariance matrix that preserves the determinant would thus preserve the probability mass within it, or its generalised variance, while potentially changing the directions, and possibly also lengths, of the semi-axes. This seems to be the closest analogue to the simple intuitive notion that can be formulated without reference to a specific choice of variables used to represent the probability distribution.
@@ 5. Parametrisations of the sub-manifold of the correlation stress tests
$g(X,Y)=$ $\frac{1}{2}\operatorname{Tr}\left(\Sigma^{-1}\frac{\partial\Sigma_{X}}{\partial t}\Sigma^{-1}\frac{\partial\Sigma_{Y}}{\partial t}\right)$ (19)
@@
As $\Sigma^{\frac{1}{2}}$ is the matrix of change of basis into one in which the covariance matrix is the identity matrix, one can interpret the above parametrisation as referring, implicitly, to changing the correlations away from the identity matrix, which is its generic form of correlations when expressed in the basis of principal components.
@@ 6. Plausibility of correlation stress tests
The line element defined by Eq. 3 is the expected value of the change of log-probability. Thus, the geodesic distance $d$ between two covariance matrices may be thought of as the integral of the infinitesimal differences between their log-probabilities, so Eq. 5 defines a natural notion of the plausibility of a stress test that morphs $\Sigma_{1}$ into $\Sigma_{2}$:
$P\left(\Sigma_{1}\rightarrow\Sigma_{2}\right)=\exp{(-d)}$ (22)
@@
$d^{2}\left(\Sigma(0),\Sigma(t)\right)=\frac{t^{2}}{2}\sum_{i=1}^{n}x_{i}^{2}$ (25)
@@
More plausible stress tests would perhaps be higher concern to a risk manager, although it is not apparent at what absolute level of of plausibility such concerns should become acute. Specifying such a level, between zero and one, is akin to explicitly specifying a level of risk tolerance to stress tests. In addition, Eq. 26 can be used to assess relative plausibility of one stress test compared to another. For example, a stress test that is tenfold, or hundredfold, as plausible as another stress test, may perhaps be considered as more relevant. Note that the plausibility of such stress tests does not depend on the initial covariance matrix as $\Sigma$ does not enter Eq. 26. Thus, the plausibility of a correlation stress test is in a sense universal: different covariance matrices can be subject to the same correlation stress test of a given plausibility.
@@ 8. Stressing along the diagonal
Of course a traceless matrix $X$ needs not have all zero diagonal elements. Without loss of generality, a convenient choice of a basis of the stress tests within the diagonal can be parametrised by setting $X_{11}=1$ and $X_{22}=-1$, with all other entries being 00:
@@ 9. Stressing all correlations of a single variable
[…]
@@ 10. Equally stressing all correlations
where the remaining $p-k$ smaller eigenvalues are all assumed to be equal to $\lambda$. This equation has a notable feature: if $p=k$, the trace of the sample covariance matrix is equal to the trace of the population covariance matrix, up to order of $O\left(m^{-2}\right)$.
@@ 11. Lax pair stress tests
Instead of the approach to correlation stress tests discussed in Section 3, one could require a different, perhaps more stringent, definition of a correlation stress test — namely that the eigenvalues of the covariance matrix do not change under such a stress test. Suppose, for simplicity, that the eigenvalues are distinct, so that the spectrum of $\Sigma(0)$ is non-degenerate. The equality of eigenvalues along a path $\Sigma(t)$ would then mean that both $\Sigma(t)$ and $\Sigma(0)$, when diagonalised, are equal:
@@
where the square brackets denote the commutator of matrices, $[A,B]=AB-BA$.
@@ 12. Example: equities, bonds, and commodities
The values of $x$, $y$, and $z$ that minimise the distance defined in Eq. 58 are 24.76, 23.84, and 649.90, respectively. Thus, the most plausible result of stressing the bond–equity correlation from 0 in $\Sigma_{1}$ to 0.1 is
@@
Figure 1: Path of bond–equity-commodity correlation stress tests

### v2

@@ 1. Introduction
This note is organised as follows: the next section introduces certain well-known features of the Fisher–Rao manifold. Section 3 discusses the meaning of covariance and correlations stress tests and Section 4 discusses the interpretation of correlation stress tests as geodesics in the Fisher–Rao manifold. Section 5 proposes an explicit parametrisation of the correlation stress tests which is the exponential map of the tangent space at the un-stressed covariance. Section 6 discusses the plausibility of the stress tests, while sections 7 through 9 discuss the stress tests that result from specific choices of “directions” in which the correlation matrix is stressed. In Section 10 we show that a correlation stress test which affects all correlations by the same amount results in the same lowest-order functional form as the estimation bias of eigenvalues of the covariance matrix due to finite amount of data. Section 11 discusses uniform sampling of correlation stress tests and introduces an algorithm for creating such samples. Section 12 reflects on certain features of the changes in the covariance matrix defined by a Lax-pair evolution and show that they would be less suitable as a model for correlation stress tests. Section 13 illustrates the main concepts using a stylised example of bonds, equities and commodities. The results are summarised in Section 14.
@@ 2. The Fisher–Rao manifold
$g_{ij}(\theta)=-\int_{\mathbb{R}^{n}}{\frac{\partial^{2}\log p(X;\theta)}{\partial\theta_{i}\partial\theta_{j}}}p(X;\theta)dX.$ (2)
@@
$d^{2}=\int_{\mathbb{R}^{n}}\left(d\log{p(X;\theta)}\right)^{2}p(X;\theta)dX.$ (3)
More specifically, consider the family of multivariate Gaussian distributions with vector of means $\mu$ and covariance $\Sigma$, and assume that both $\mu$ and $\Sigma$ are differentiable functions of some arbitrary parameters $\theta_{i}$, $i=1,\dots,k$ . The Fisher–Rao metric reduces to [3, 18]:
@@ 3. Covariance and correlation stress tests
A geometric interpretation provides some further guidance: the determinant of the covariance matrix is, for a multivariate normal distribution, the volume of an ellipsoid of constant probability density enveloping a given probability mass. The directions of the semi-axes of this ellipsoid represent the uncorrelated directions (principal components), while the length of a semi-axis represents the standard deviation of the principal component. A change in the covariance matrix that preserves the determinant would thus preserve the probability mass within it, or its generalised variance, while potentially changing the directions, and possibly also the lengths, of the semi-axes. This seems to be the closest analogue to the simple intuitive notion that can be formulated without reference to a specific choice of variables used to represent the probability distribution.
@@ 5. Parametrisations of the sub-manifold of the correlation stress tests
$g(X,Y)=$ $\frac{1}{2}\operatorname{Tr}\left(\Sigma^{-1}\frac{\partial\Sigma_{X}}{\partial t}\Sigma^{-1}\frac{\partial\Sigma_{Y}}{\partial t}\right)$
@@
As $\Sigma^{\frac{1}{2}}$ is the matrix of change of basis into one in which the covariance matrix is the identity matrix, one can interpret the above parametrisation as implicitly referring to changing the correlations away from the identity matrix, when expressed in the basis of the normalised eigenvectors of $\Sigma$.
@@ 6. Plausibility of correlation stress tests
The line element defined by Eq. 3 is the expected value of the change of the log-probability corresponding to an infinitesimal change of the distribution. Intuitively, the more a distribution is changed, the less plausible the new distribution may be. The geodesic distance $d$ between two distributions is a natural measure of how much these two distributions differ. It may be thought of as the integral of the infinitesimal differences between their logarithms, hence the exponential function of that distance can be interpreted as a measure of the plausibility of a stress test that morphs a distribution $p_{1}(X)$ into another distribution $p_{2}(X)$,
$P\left(p_{1}\rightarrow p_{2}\right)=\exp{(-d)}.$ (21)
The plausibility will range between zero and one, the former corresponding to an infinitely large stress, and the latter to no stress at all.
For multivariate normal distributions, the plausibility of a stress test that morphs $\Sigma_{1}$ into $\Sigma_{2}$ is
$P\left(\Sigma_{1}\rightarrow\Sigma_{2}\right)=\exp{(-d)}.$ (22)
@@
$d^{2}\left(\Sigma(0),\Sigma(t)\right)=\frac{t^{2}}{2}\sum_{i=1}^{n}x_{i}^{2}=\frac{t^{2}}{2}\operatorname{Tr}{\left(X^{2}\right)}$ (25)
@@
More plausible stress tests would perhaps be of higher concern to a risk manager. Specifying a plausibility threshold level, between zero and one, is akin to explicitly specifying a level of risk tolerance to stress tests. One can also use Eqs. 21 and 26 to assess relative plausibility of one stress test compared to another. For example, a stress test that is tenfold, or hundredfold, as plausible as another stress test, may perhaps be considered as more relevant.
Note that the plausibility of a correlation stress test does not depend on the initial covariance matrix as $\Sigma(0)$ does not enter Eq. 26. Thus, the plausibility of a correlation stress test is in a sense universal: different covariance matrices can be subjected to the same correlation stress test of a given plausibility, and stress tests can be ranked according to their plausibility irrespective of the initial unstressed covariance matrix.
@@ 8. Stressing along the diagonal
Of course a traceless matrix $X$ needs not have all zero diagonal elements. Without loss of generality, a convenient choice of a basis of the stress tests within the diagonal can be parametrised by setting $X_{11}=1$ and $X_{22}=-1$, with all other entries being 0:
@@ 9. Stressing all correlations of a single variable
$X^{k}=\begin{dcases*}(n-1)^{\frac{k}{2}-1}Z&when $k$ is even\\ (n-1)^{\frac{k-1}{2}}X&when $k$ is odd.\end{dcases*}$ (37)
@@ 10. Equally stressing all correlations
where the remaining pkp-k smaller eigenvalues are all assumed to be equal to $\lambda$. This equation has a notable feature: if $p=k$, the trace of the sample covariance matrix is equal to the trace of the population covariance matrix, up to order of $O\left(m^{-2}\right)$.
@@
11. Uniform sampling of correlation stress tests
Eqs. 15 and 26 show how to define an arbitrary correlation stress test at a given plausibility level $P$. In practice, it may also be useful to be sample such stress tests in a way that does not rely on a specific preference of one such a specific stress over another. Thus, one would need to “uniformly” sample correlation stress tests. Uniformity is closely related to measures ad symmetries: a sample is uniformly distributed if it has the same number of points within sets of the same measure, and such a measure needs to be invariant under the symmetries of the problem.
Here, the set of points is selected from the stressed covariances with the same plausibility, which are the points at a fixed geodesic distance from the unstressed matrix $\Sigma_{0}$. These points form a normal sphere — the submanifold of a given Rao distance from the unstressed matrix. The symmetry of this submanifold is directly related to the symmetry of the tangent space at $\Sigma_{0}$ which preserves the Frobenius scalar product, Eq. 19. This can be shown as follows: select an arbitrary point on the normal sphere, and consider its pull back to the tangent space at $\Sigma_{0}$. Then, rotate the resulting tangent vector through a linear transformation which is orthogonal with respect to the Frobenius norm. Finally, push forward the so-rotated tangent vector to the normal sphere  [^1] This would normally be true only up to a certain value of the radius of the normal sphere. The Fisher–Rao manifold considered here has negative curvature however, so is true for any radius. Thus, the group symmetries of the normal sphere is the same as the group of symmetries of the scalar-product preserving symmetries of the tangent space, and both groups act transitively on their respective domains. As a result, there is a unique up to a multiplicative constant measure that is invariant under these groups, the Haar measure. Moreover, both the group of transformation of the normal sphere and the group of transformations on the tangent space act transitively. Thus the measure on the normal sphere and the measure on the sphere in the tangent space are the same, up to a multiplicative constant, namely the standard measure on a sphere in the tangent space.
The tangent space at $\Sigma_{0}$ consists of symmetric traceless matrices. The Frobenius norm can be expressed in the standard vector notation as
$\vec{X}\cdot\vec{Y}=\operatorname{Tr}{XY},$ (51)
where $X$ and $Y$ are general $n\times n$ matrices viewed as vectors in $\mathrm{R}^{N}$, where $N=n^{2}$, and $n$ is number of risk factors. The unit sphere in this space is the set of points such that
$\vec{X}\cdot\vec{X}=\operatorname{Tr}{XX}=1.$ (52)
Eq. 52 defines the unit sphere in the $N$-dimensional space of general nnxnn matrices. Symmetric and trace-free matrices form a subspace thereof which is an orthogonal projection of $\mathrm{R}^{N}$ with respect to a set of mutually orthogonal vectors. Thus, the projection of the unit sphere in $\mathrm{R}^{N}$ is also a unit sphere of lower dimension, n(n1)/21n(n-1)/2-1. This can be shown as follows.
Let $E_{ab}$, with $a,b\in 1,\dots,n$ ,be matrices whose entries are all zero except for the entry, in $a$-th row and $b$-th column, which is $1$ and the entry in $b$-th row and $a$-th column, which is set to 1-1. Let $I$ be the identity matrix. Since
$\vec{E}_{ab}\cdot\vec{E}_{cd}=\operatorname{Tr}{E_{ab}E_{cd}}=0,$ (53)
and
$\vec{E}_{ab}\cdot\vec{I}=\operatorname{Tr}{E_{ab}I}=0,$ (54)
the matrices $E_{ab}$ and $I$ span a set of mutually orthogonal vectors in vector space $R^{N}$ of general $n\times n$ matrices.
Suppose that a matrix $X$ satisfies
$\vec{E}_{ab}\cdot\vec{X}=0,$ (55)
for any $a,b\in 1,\dots,n$ and
$\vec{I}\cdot\vec{X}=\operatorname{Tr}{IX}=0.$ (56)
Then $X$ is trivially symmetric and trace-free. As the $N-1=n^{2}-1$ dimensional sphere consisting of all matrices $X$ such that
$\vec{X}\cdot\vec{X}=\operatorname{Tr}{XX}=1$ (57)
is (spherically) symmetric, its projection onto the subspace orthogonal to all matrices $E_{ab}$ and $I$ will thus be also a sphere, in the natural metric induced from $R^{N}$ and thus provides the notion of uniformity for symmetric positive definite matrices.
The arguments presented above suggest the following algorithm for constructing a uniformly distributed set of $M$ correlation stress tests $\Sigma_{s}$ of an unstressed covariance matrix $\Sigma_{0}$:
  1. 1.
    Select $N^{2}$ independently and normally distributed random numbers. Rescale them so that the sum of their squares is $1$. The rotational symmetry of the uncorrelated multivariate normal distribution implies that such numbers, when viewed as coordinates in $\mathrm{R}^{N}$, represent a matrix $X$ from a population that is uniformly distributed on a unit sphere S(N1)S(N-1).
  2. 2.
    For all $a,b\in 1,\dots,n$ with $a<b$, recursively project this matrix onto the sub-space orthogonal to each consecutive $E_{ab}$,
    $\vec{X}\coloneqq\vec{X}-\frac{\vec{X}\cdot\vec{E}_{ab}}{\vec{E}_{ab}\cdot\vec{E}_{ab}}E_{ab},$ (58)
    where $\coloneqq$ denotes the assignment for the next recursion step.
  3. 3.
    Project this matrix further onto the sub-space orthogonal to the unit matrix $I$:
    $\vec{X}\coloneqq\vec{X}-\frac{\vec{X}\cdot\vec{I}}{n}\vec{I}.$ (59)
  4. 4.
    Finally, use Eq. 15
    $\Sigma_{s}\coloneqq\Sigma_{0}^{\frac{1}{2}}\exp{(Xt)}\Sigma_{0}^{\frac{1}{2}}$ (60)
    to define the stressed covariance matrix which is a point on the normal sphere of radius defined by Eq. 25
Repeating the above steps $M$ times will result in a sample of uniformly distributed covariance matrices, each corresponding to a correlation stress of the same plausibility, as defined in Eq. 26.
An example of applying the above algorithm to a stylised case of portfolio of three assets is discussed in Section 13.
@@ 12. Lax pair stress tests
Instead of the approach to correlation stress tests discussed in Section 3, one could require a different, perhaps more stringent, definition of a correlation stress test — namely that the eigenvalues of the covariance matrix do not change under such a stress test. In this section, we show that such an approach would be less suitable than that of Eq. 15.
Suppose, for simplicity, that the eigenvalues are distinct, so that the spectrum of $\Sigma(0)$ is non-degenerate. The equality of eigenvalues along a path $\Sigma(t)$ would then mean that both $\Sigma(t)$ and $\Sigma(0)$, when diagonalised, are equal:
@@
where the square brackets denote the commutator of matrices, [A,B]=ABBA[A,B]=AB-BA.
@@ 13. Example: equities, bonds, and commodities
The values of $x$, $y$, and $z$ that minimise the distance defined in Eq. 68 are 24.76, 23.84, and 649.90, respectively  [^2] The manifold of positive-definite symmetric matrices has non-positive sectional curvature, and the extrinsic curvature of the sub-manifold of matrices with a given determinant vanishes [1]. Thus, by the Gauss equation, the submanifold has non-positive sectional curvature, which through the Cartan-Hadamard theorem implies that the minimum is global.. Thus, the most plausible result of stressing the bond–equity correlation from 0 in $\Sigma_{1}$ to 0.1 is
@@
Figure 1: Path of bond–equity-commodity correlation stress tests. The solid line corresponds to the matrix of Eq. 71 while the dashed line corresponds to the stress with plausibility of 0.49.
@@
We will now consider a random sample, of size $M=10000$, of uniformly distributed stressed covariance matrices with plausibility $\exp(-1/\sqrt{2})\approx 0.49$ with respect to the covariance matrix of Eq. 69. This sample is obtained using the algorithm described in Section 11. Figure 2 shows the distribution of the effective number of degrees of freedom – or the exponential of the Shannon entropy of the eigenvalues of $\lambda_{i}$ of these matrices:
$N_{df}=\exp{\left(-\sum_{i=1}^{N}p_{i}\log{p_{i}}\right)},$ (73)
where
$p_{i}=\left(\sum_{i=1}^{N}\lambda_{i}\right)^{-1}\lambda_{i}.$ (74)
Figure 3 shows the distribution of the ex-ante volatility, calculated using these matrices, for a pro-forma portfolio of 50% equities, 40% bonds, and 10% commodities. One can observe that the ex-ante volatility for a meaningful fraction of such matrices is substantially higher than that of either the unstressed, or the fitted, covariance matrix. The effect would be of course more pronounced for stress tests with lower, but still meaningful, plausibility.
Figure 2: Number of independent bets for a sample of plausibility 0.49 (indicated by the dashed line in Fig. 1). The black vertical lines correspond to 1.87, the number of bets for the fitted matrix given by Eq.71, and 1.90, for the unstressed covariance matrix.
Figure 3: Ex-ante volatility for a sample of plausibility 0.49. The black vertical lines correspond to the volatility of 6.92% for the unstressed covariance matrix and 7.45% for the fitted matrix of Eq. 71